# Rethinking Softmax: Self-Attention with Polynomial Activations

### v1

@@ Abstract
This paper challenges the conventional belief that softmax attention in transformers is effective primarily because it generates a probability distribution for attention allocation. Instead, we theoretically show that its success lies in its ability to implicitly regularize the Frobenius norm of the attention matrix during training. We then explore alternative activations that regularize the Frobenius norm of the attention matrix, demonstrating that certain polynomial activations can achieve this effect, making them suitable for attention-based architectures. Empirical results indicate these activations perform comparably or better than softmax across various computer vision and language tasks, suggesting new possibilities for attention mechanisms beyond softmax.
@@ 1 Introduction
Transformer architectures (Vaswani et al., 2017) have become the state-of-the-art model architecture for diverse areas such as natural language processing (NLP) (Vaswani et al., 2017; Devlin et al., 2018; Zhuang et al., 2021; Zhen et al., 2022), computer vision (Dosovitskiy et al., 2020; Carion et al., 2020; Liu et al., 2021; Touvron et al., 2021), and robotics (Fu et al., 2024; Maiti et al., 2023; Salzmann et al., 2020). A key component in the transformer architecture is the softmax attention block, enabling transformers to evaluate the importance of individual input elements during output generation. This feature facilitates an efficient method to attend to diverse input elements throughout training, allowing transformers to effectively capture spatial dependencies within sequential data. Unlike traditional recurrent neural networks (RNNs) and convolutional neural networks (CNNs), transformers exhibit the ability to scale to large datasets without a significant degradation in performance. This characteristic has made them an ideal architecture for handling large-scale machine learning tasks.
Softmax is widely recognized for its effectiveness in attention mechanisms due to its ability to produce an attention matrix that meets three key conditions (Vaswani et al., 2017; Dosovitskiy et al., 2020; Zhen et al., 2022; AUEB et al., 2016): (i) non-negativity, (ii) rows that are normalized to sum to 1 and (iii) sparsity. The general consensus is that non-negativity guarantees that attention weights remain positive, facilitating the model in assigning significance to various input elements. The normalization constraint ensures that the attention weights for all input elements collectively sum to $1$, rendering the weights interpretable as probabilities. Additionally, sparsity aids the model in focusing on a select few key elements, thereby enhancing efficiency. It has been argued that these properties are crucial for enabling the attention mechanism to attend to pertinent segments of the input sequence while efficiently filtering out irrelevant details. However, this approach to attention has become somewhat axiomatic as it is mostly motivated by empirical results with little theoretical foundation. Despite the exploration of alternative activations in several studies (Shen et al., 2023; Fang et al., 2022; Correia et al., 2019), softmax attention continues to dominate, largely due to its interpretability.
In this paper, we question this view by proposing that the effectiveness of softmax stems from its implicit regularization of the Frobenius norm of the attention matrix during training, preventing attention weights from becoming excessively large or small. We then derive a theoretical framework that produces polynomial activations that deliberately violate one or more of the three conditions mentioned earlier, yet are able to regularize the Frobenius norm of the attention weights during training. Our findings demonstrate that such activations can achieve comparable or even superior performance to softmax across various vision and natural language processing (NLP) tasks, even though they seem to violate our understanding of attention.
We advise the reader that this paper diverges from the usual pursuit of creating cutting-edge transformer architectures for achieving state-of-the-art results on benchmark datasets. Instead, our focus is on critically examining softmax attention to determine whether its effectiveness is a result of true interpretability or a more nuanced regularization mechanism. By questioning established views, we aim to uncover deeper insights into transformer architectures that could lead to broader applications and improved understanding. Nonetheless, we validate our theory on multiple transformer-based tasks, including image classification, segmentation, object detection, and NLP, often achieving results that match or surpass those of softmax attention. Our main contributions are:
@@
    We question the widely accepted notion that softmax’s effectiveness in attention mechanisms is solely due to its ability to produce normalized sparse attention weights. Instead, we theoretically show that softmax has a regularization effect on attention and argue this plays a crucial role in its success.
@@
    We explore activations that deliberately deviate from traditional softmax attention conditions. These activations are found to regularize the Frobenius norm of the attention matrix during training, akin to softmax, and demonstrate comparable or superior performance across various vision and NLP tasks.
@@ 2 Related Work
Several studies have explored alternative activations for attention mechanisms in transformers. Shen et al. (2023) investigated ReLU activations, finding them to outperform softmax in tasks with long sequences, such as document translation. Banerjee et al. (2020) examined Taylor series approximations to softmax, which showed superior performance to softmax in image classification. Wang et al. (2021) proposed periodic alternatives to softmax, designed to provide better gradients for attention mechanisms and achieved better results than softmax on simple networks for image classification. Koohpayegani & Pirsiavash (2024) demonstrated that applying $l^{1}$ normalization to linear attention mechanisms can achieve performance comparable to softmax. Our work differs from all these in that we identify a clear theoretical relationship between the scale of the Frobenius norm of the self-attention matrix and the input sequence length. Using this insight to derive potential acitvations that can perform on par with softmax.
@@ 3 Preliminaries and Notation
In this section we outline the definition of a transformer via the transformer block and set the notation of various mathematical quantities we will be using in future sections. For more details on transformers the reader can consult Vaswani et al. (2017); Dosovitskiy et al. (2020).
@@
$\mathbf{T}(x)=\mathbf{F}(\mathbf{A}(x)+x)$ (3.1)
@@
$\mathbf{A}(X)=\mathbf{softmax}\bigg{(}\frac{qk^{T}}{\sqrt{d}}\bigg{)}v=\mathbf{softmax}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}XV.$ (3.3)
@@
$\mathbf{softmax}\bigg{(}\begin{bmatrix}x_{11}&\cdots&x_{1n}\\ \vdots&\vdots&\vdots\\ x_{n1}&\cdots&x_{nn}\end{bmatrix}\bigg{)}=\begin{bmatrix}\frac{e^{x_{11}}}{\sum_{j=1}^{n}e^{x_{1j}}}&\cdots&\frac{e^{x_{1n}}}{\sum_{j=1}^{n}e^{x_{1j}}}\\ \vdots&\vdots&\vdots\\ \frac{e^{x_{n1}}}{\sum_{j=1}^{n}e^{x_{nj}}}&\cdots&\frac{e^{x_{nn}}}{\sum_{j=1}^{n}e^{x_{nj}}}\end{bmatrix}$ (3.4)
The factor $\frac{1}{\sqrt{d}}$, as explained in Vaswani et al. (2017), is a scaling to prevent the gradients of softmax from being too small. For the theoretical analysis in this paper we will only use the dot-product similarity $qk^{T}$ and call the $N\times N$ matrix $softmax(qk^{T})$ the softmax self-attention matrix. In the experiments section, Sec. 5, we will empirically validate our theoretical framework on more general softmax attention blocks.
For general transformer architectures, multiple heads $\mathbf{A}_{i}$ for $1\leq i\leq n$ are used. Each attention head is defined by equation 3.3 and then all outputs of each attention head are concatenated together before going into the feedforward layer.
We will need notation for the derivative of the matrix softmax map defined by equation 3.4. Given a matrix $A\in\mathbb{R}^{N\times N}$ we can differentiate the matrix map $\mathbf{softmax}$ at $A$ and obtain the gradient linear map $\mathbf{\nabla softmax}(A):\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ that is defined by the formula
$\mathbf{\nabla softmax}(A):=\mathbf{Jsoftmax}(A)^{T}$ (3.5)
@@ 4.1 Implicit regulatization of Softmax
This section presents a theoretical result showing that the softmax activation imposes control over the Frobenius norm of the self-attention matrix in a way that grows sub-linearly with the input sequence’s token length. Additionally, we demonstrate that the gradient of the softmax with respect to the self-attention matrix also exhibits a similar degree of regularity. While previous work has analyzed the regularity of softmax self-attention through the lens of the Lipschitz constant (Kim et al., 2021; Castin et al., 2023), our theorem offers a novel perspective by directly linking the Frobenius norm regularity to the token length. This provides insights into how self-attention activations should scale with token length to maintain stability during training, especially with gradient descent-based algorithms.
@@
Let $\mathbf{softmax}:\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ be the matrix softmax map defined by equation 3.4 and let $\mathbf{\nabla softmax}(A):\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ denote the gradient of $\mathbf{softmax}$ at $A\in\mathbb{R}^{N\times N}$. We then have the following bounds on the Frobenius norms
$||\mathbf{softmax}(A)||_{F}$ $\leq\sqrt{N}$ (4.1)
$||\mathbf{\nabla softmax}(A)||_{F}$ $\leq 2\sqrt{N}.$ (4.2)
The key implication of theorem 4.1 is that during the training of a transformer with softmax self-attention, the Frobenius norm of each softmax self-attention matrix remains bounded by a value that grows as $\mathcal{O}(\sqrt{N})$. This ensures that backpropagation through the weights of the self-attention matrix does not lead to excessively large gradients. The proof hinges on the fact that the row normalization inherent in softmax effectively controls the Frobenius norm. For a detailed proof see appendix A.1.1.
@@ 4.2 Polynomial activations for self-attention
$\mathbb{E}\bigg{|}\bigg{|}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(N)$ (4.3)
@@
$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(\sqrt{N}).$ (4.4)
Corollary 4.3 establishes that activations of the form $\phi(x):=\frac{1}{\sqrt{N}}x^{p}$ provide a level of regularization, in expectation, similar to that of softmax when applied to the self-attention matrix. The proof of theorem 4.2 can be found in appendix A.1.2. The next property we want to prove is one similar to the gradient bound obtained in theorem 4.1. Since the self-attention matrix has parameters given by the queries $Q$ and keys $K$ (Vaswani et al., 2017), this implies that during the training of a transformer the $Q$ and $K$ matrices are the only aspects of the self-attention matrix that get updated. Therefore, we compute a regularity result with respect to the $Q$ and $K$ derivatives.
@@
$\mathbb{E}\bigg{|}\bigg{|}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(N)$ (4.5)
The above theorem then suggests that if we scale the polynomial $x\rightarrow x^{p}$ by $\frac{1}{\sqrt{N}}$ the $Q$ derivative will have $\mathcal{O}(\sqrt{N})$ growth.
@@
$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(\sqrt{N}).$ (4.6)
@@
Corollaries 4.3 and 4.5 suggest that polynomial activations of the form $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$, with $p>0$, can achieve performance comparable to softmax when applied to self-attention matrices. In section 5, we empirically compare these activations to softmax and observe that they outperform softmax on a variety of transformer tasks. We focus on $p=1$ and $p=3$ as these polynomials clearly violate key aspects of softmax based attention, such as normalized rows, positivity, and sparsity. For larger values of $p$, performance declines due to the functions $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$ having smaller gradients around 00 when $p$ is large, causing difficulties in training.

### v2

# Rethinking Attention: Polynomial Alternatives to Softmax in Transformers

@@ Abstract
This paper questions whether the strong performance of softmax attention in transformers stems from producing a probability distribution over inputs. Instead, we argue that softmax’s effectiveness lies in its implicit regularization of the Frobenius norm of the attention matrix, which stabilizes training. Motivated by this, we explore alternative activations, specifically polynomials, that achieve a similar regularization effect. Our theoretical analysis shows that certain polynomials can serve as effective substitutes for softmax, achieving strong performance across transformer applications despite violating softmax’s typical properties of positivity, normalization, and sparsity. Extensive experiments support these findings, offering a new perspective on attention mechanisms.
@@ 1 Introduction
Transformer architectures [26] have become the foundation of state-of-the-art models across natural language processing (NLP) [26, 9, 33, 32], computer vision [10, 6, 17, 25], and robotics [12, 18, 19]. A core component of these models is the softmax attention block, which computes token interactions by evaluating the relative importance of inputs. This mechanism enables transformers to effectively model long-range dependencies, a capability that distinguishes them from recurrent neural networks (RNNs) and convolutional neural networks (CNNs), particularly in scaling to large datasets.
Softmax self-attention satisfies three key properties: (1) non-negativity of attention weights, (2) row-wise normalization to ensure weights sum to one (interpretable as probabilities), and (3) sparsity, promoting focus on a few relevant tokens. These properties are commonly believed to be essential for modeling effective attention [4, 32], facilitating interpretability. However, this viewpoint has largely been adopted on empirical grounds, with limited theoretical justification. Despite investigations into alternative activations [20, 11, 8], softmax remains dominant, primarily due to its empirical performance and interpretability.
In this paper, we challenge this prevailing assumption and ask:
{greenbox}
Do attention mechanisms in transformers require non-negativity, normalization, and sparsity for effective performance?
We present a new perspective suggesting that softmax’s effectiveness stems not inherently from these properties, but from its implicit regularization of the attention matrix’s Frobenius norm during training. Building on this insight, we show that simple polynomial activations, which violate non-negativity, normalization, and sparsity, can nevertheless provide comparable or superior performance to softmax by inducing a similar regularization effect. This offers a fundamentally different interpretation of attention, distinct from the conventional probabilistic view of softmax.
Our contributions are as follows:
@@
    We provide a theoretical analysis showing that softmax attention implicitly regularizes the Frobenius norm of the attention matrix, challenging the assumption that non-negativity, normalization, and sparsity are the primary drivers of its success.
@@
    We propose polynomial activations as an alternative to softmax, demonstrating that they can induce similar regularization effects without adhering to the traditional softmax constraints, achieving competitive performance across tasks including image classification, object detection, instance segmentation, text classification, and physics-based modeling.
By revisiting foundational assumptions, our work offers deeper insights into the mechanics of attention and opens new avenues for designing alternative attention mechanisms.
@@ 2 Related Work
Attention activations.
A variety of alternative activations for attention mechanisms have been explored in recent literature. [20] showed that ReLU activations outperform softmax in long-sequence tasks, such as document translation. Of particular relevance to our work, [29] demonstrated that scaling ReLU by the inverse of sequence length can surpass softmax in certain vision applications, emphasizing the importance of correct activation scaling. In this paper, we show that softmax inherently applies such a scale through its normalization, and we derive theoretical principles that motivate polynomial activations with scalings proportional to the square root of the sequence length. Other studies have proposed alternatives with varying motivations. [5] used Taylor series approximations of softmax, achieving superior performance in image classification. [27] introduced periodic activations to improve gradient flow in attention layers. [15] showed that $l^{1}$ normalization applied to linear attention mechanisms can yield on par performance to softmax’s on three distinct vision transformers. Distinct from these works, our approach establishes a clear theoretical link between the Frobenius norm of the attention matrix and the input sequence length. Leveraging this insight, we design polynomial activations that break three canonical properties of softmax—non-negativity, row normalization, and sparsity—while still achieving competitive performance.
Attention mechanisms.
Numerous strategies have been proposed to improve the efficiency and scalability of transformers by reducing computational overhead and rethinking attention mechanisms. The Data-Efficient Image Transformer (DeiT) [25] leverages distillation tokens to achieve competitive performance without relying on large datasets. The Cross-Covariance Image Transformer (XCiT) [3] introduces cross-covariance attention, enabling efficient spatial interactions with reduced complexity. The Swin Transformer [17] employs a hierarchical architecture with shifted window-based self-attention to enhance scalability for vision tasks. The Nyströmformer [30] approximates full self-attention using the Nyström method, reducing its complexity from quadratic to near-linear. Similarly, the MLP-Mixer [24] replaces self-attention entirely with multi-layer perceptrons for spatial and channel mixing. In this work, we demonstrate that our polynomial-based attention activations can be integrated into these architectures [25, 3, 17, 30], matching softmax performance while violating its typical properties of non-negativity, normalization, and sparsity.
@@ 3 Preliminaries and Notation
In this section we outline the definition of a transformer via the transformer block and set the notation of various mathematical quantities we will be using in future sections. For more details on transformers the reader can consult [26, 10].
@@
$\mathbf{T}(x)=\mathbf{F}(\mathbf{A}(x)+x)$ (1)
@@
$\begin{split}\mathbf{A}(X)&=\mathbf{softmax}\bigg{(}\frac{qk^{T}}{\sqrt{d}}\bigg{)}v\\ &=\mathbf{softmax}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}XV.\end{split}$ (3)
@@
$\mathbf{softmax}\bigg{(}\begin{bmatrix}x_{11}&\cdots&x_{1n}\\ \vdots&\vdots&\vdots\\ x_{n1}&\cdots&x_{nn}\end{bmatrix}\bigg{)}=\begin{bmatrix}\frac{e^{x_{11}}}{\sum_{j=1}^{n}e^{x_{1j}}}&\cdots&\frac{e^{x_{1n}}}{\sum_{j=1}^{n}e^{x_{1j}}}\\ \vdots&\vdots&\vdots\\ \frac{e^{x_{n1}}}{\sum_{j=1}^{n}e^{x_{nj}}}&\cdots&\frac{e^{x_{nn}}}{\sum_{j=1}^{n}e^{x_{nj}}}\end{bmatrix}$ (4)
The factor $\frac{1}{\sqrt{d}}$, as explained in [26], is a scaling to prevent the gradients of softmax from being too small. For the theoretical analysis in this paper we will only use the dot-product similarity $qk^{T}$ and call the $N\times N$ matrix $softmax(qk^{T})$ the softmax self-attention matrix. In the experiments, Section 6, we will empirically validate our theoretical framework on more general softmax attention blocks used in state of the art transformers such as DeiT [25], Swin Transformer [17] and XciT [30].
For general transformer architectures, multiple heads $\mathbf{A}_{i}$ for $1\leq i\leq n$ are used. Each attention head is defined by (3) and then all outputs of each attention head are concatenated together before going into the feedforward layer.
We will need notation for the derivative of the matrix softmax map defined by (4). Given a matrix $A\in\mathbb{R}^{N\times N}$ we can differentiate the matrix map $\mathbf{softmax}$ at $A$ and obtain the gradient linear map $\mathbf{\nabla softmax}(A):\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ that is defined by the formula
$\mathbf{\nabla softmax}(A):=\mathbf{Jsoftmax}(A)^{T}$ (5)
@@ 4.1 Implicit regulatization of Softmax
This section presents a theoretical result showing that the softmax activation imposes control over the Frobenius norm of the self-attention matrix in a way that grows sub-linearly with the input sequence’s token length. Additionally, we demonstrate that the gradient of the softmax with respect to the self-attention matrix also exhibits a similar degree of regularity. While previous work has analyzed the regularity of softmax self-attention through the lens of the Lipschitz constant [14, 7], our theorem offers a novel perspective by directly linking the Frobenius norm regularity to the token length. This provides insights into how self-attention activations should scale with token length to maintain stability during training, especially with gradient descent-based algorithms.
{bluebox}
@@
Let $\mathbf{softmax}:\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ be the matrix softmax map defined by (4) and let $\mathbf{\nabla softmax}(A):\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ denote the gradient of $\mathbf{softmax}$ at $A\in\mathbb{R}^{N\times N}$. We then have the following bounds on the Frobenius norms
$||\mathbf{softmax}(A)||_{F}$ $\leq\sqrt{N}$ (6)
$||\mathbf{\nabla softmax}(A)||_{F}$ $\leq 2\sqrt{N}.$ (7)
The key implication of theorem 4.1 is that during the training of a transformer with softmax self-attention, the Frobenius norm of each softmax self-attention matrix remains bounded by a value that grows as $\mathcal{O}(\sqrt{N})$. This ensures that backpropagation through the weights of the self-attention matrix does not lead to excessively large gradients. The proof hinges on the fact that the row normalization inherent in softmax effectively controls the Frobenius norm. For a detailed proof see Section A.1.1.
@@ 4.2 Polynomial activations for self-attention
$\mathbb{E}\bigg{|}\bigg{|}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(N)$ (8)
@@
$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(\sqrt{N}).$ (9)
Corollary 4.3 establishes that activations of the form $\phi(x):=\frac{1}{\sqrt{N}}x^{p}$ provide a level of regularization, in expectation, similar to that of softmax when applied to the self-attention matrix. The proof of theorem 4.2 can be found in appendix A.1.2. The next property we want to prove is one similar to the gradient bound obtained in Theorem 4.1. Since the self-attention matrix has parameters given by the queries $Q$ and keys $K$ [26], this implies that during the training of a transformer the $Q$ and $K$ matrices are the only aspects of the self-attention matrix that get updated. Therefore, we compute a derivative bound with respect to the $Q$ and $K$ derivatives.
{bluebox}
@@
$\mathbb{E}\bigg{|}\bigg{|}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(N)$ (10)
Figure 1: Training ViT-Tiny with the activation $\phi(x)=x^{3}$ with different sequence lengths and different scales. As the sequence length gets larger, the $k$ scale (x-axis) needed to obtain good accuracy when using $\frac{1}{k}x^{3}$ as an activation increases validating the theory from section 4.2.
{orangebox}
@@
$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(\sqrt{N}).$ (11)
@@
Remark 4.6.
Corollary 4.3 and Corollary 4.5 suggest that polynomial activations of the form $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$, with $p>0$, can achieve performance comparable to softmax when applied to self-attention matrices. We point out that both corollaries rested on the assumption that $X$, $Q$ and $K$ are i.i.d random variables. In general, this is only true at initialization when training a transformer [2]. Although this is a limitation in the theory the experiments in Section 6 show that this insight can be used effectively to develop new attention blocks that perform comparable to softmax yet violate the three conditions of positivity, normalized rows summing to $1$ and sparsity showing that attention blocks do not need to be modeled as a probability distribution.
5 Testing the theory
In this section we test the theory developed in Section 4.2 on small vision transformers. We will consider the activation $\phi(x)=\frac{1}{k}x^{3}$, where $k>0$ is a fixed scale, as this activation clearly violates the three key conditions of softmax based attention; positivity, normalization and sparsity. We found that polynomials $\frac{1}{k}x^{p}$ for $p>3$ did not perform well during training as they witnessed a gradient vanishing problem due to the fact that the function $x^{p}$ for $p$ large have very small values around 00.
The first experiment we conducted was to test how the Top-$1\%$ accuracy changes for a ViT-Tiny vision transformer [22], trained on Tiny-Imagenet dataset [21], as we change the sequence length $N$ of the input and the scale predicted in corollaries 4.3 and 4.5 when using the activation $\frac{1}{k}x^{3}$. The standard ViT-Tiny model comprises $12$ transformer layers, each equipped with $3$ attention heads, with each head having dimension $64$. We considered four different input sequence lengths $N$ of sizes $256$, $64$, $16$ and $8$. For each such sequence length, we ran a ViT-Tiny architecture with the activation $\frac{1}{k}x^{3}$ where $k$ ranged from roughly $10^{-3}$ to $10^{3}$. According to the theory developed in section 4, the Frobenius norm of $\frac{1}{\sqrt{N}}x^{3}$ scales according to $\mathcal{O}(\sqrt{N})$. Thus the best accuracy should occur when $k=\mathcal{O}(\sqrt{N})$ and should degrade for other values due to training instability. Figure 1 shows the results of the experiment, we note that the $x-axis$ plots the values of $k$. We see that as the sequence length increases the factor of $k$ needs to increase so that the activation $\frac{1}{k}x^{3}$ performs well on the ViT-Tiny architecture as predicted by the theory developed in corollaries 4.3 and 4.5.
In a second experiment we decided to compare the performance of the original ViT-Tiny architecture, that uses an input sequence length of $256$, with a softmax activation and a polynomial activation on the Tiny-ImageNet dataset. In Corollary 4.3 and Corollary 4.5 it was pointed out that the scaling of the polynomial is important to keep the Frobenius norm of the polynomial attention matrix from becoming too large. The results of those corollaries suggested that a scale to use is $\frac{1}{\sqrt{N}}$, $N$ being the sequence length, which in this case is $\frac{1}{\sqrt{256}}=\frac{1}{16}$. We therefore decided to to compare the three activations softmax, $x^{3}$ and $\frac{1}{16}x^{3}$. Each was trained for 200 epochs using the AdamW optimizer. To begin with we computed the Frobenius norm of the attention matrix throughout training averaged over all the heads in layers 2, 7 and 12 of the ViT-Tiny architecture when trained on the Tiny-ImageNet dataset. Figure 2 plots the results. We see from the figure that the Frobenius norm of of the $x^{3}$ archicture is much larger than softmax but scaling it by $\frac{1}{16}$ brings it down to softmax levels. Similarly, Figure 3 shows the Jacobian’s Frobenius norm, where scaling also brings the norms closer to softmax, ensuring more stable gradients. Further plots for other layers are in Section A.2.3. Table 1 presents the final Top-1% accuracy achieved by each activation function. Notably, $\frac{x^{3}}{16}$ delivers the best performance. In contrast, the unscaled $x^{3}$ activation yields significantly lower Top-1% accuracy, underscoring the importance of incorporating an appropriate scaling factor.
Figure 2: Frobenius norm of the self-attention matrix with three different activations in layer 2, 7 and 12 of the ViT-Tiny architecture during training.
Figure 3: Frobenius norm of the jacobian of the self-attention matrix with three different activations in layer 2, 7 and 12 of the ViT-Tiny architecture during training.
Table 1: Comparison of Top-1% accuracy on Tiny-ImageNet between softmax and polynomial activations.
softmax $\frac{x^{3}}{16}$ $x^{3}$
Top-1% accuracy 50.26 50.5 45.3